# Fine-Grained Representation Learning via Multi-Level Contrastive Learning without Class Priors

### v2

@@ Abstract
Recent advances in unsupervised representation learning frequently leverage class information to improve the extraction and clustering of features. However, this dependence on class priors limits the applicability of such methods in real-world scenarios where class information is unavailable or ambiguous. In this paper, we propose Contrastive Disentangling (CD), a simple yet effective framework that learns representations without any relying on class priors. CD employs a multi-level contrastive learning strategy, integrating instance-level and feature-level losses with a normalized entropy loss to learn semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss encourages the separation of feature representations between different samples; (2) the feature-level contrastive loss promotes independence among feature prediction heads; and (3) the normalized entropy loss ensures that the feature heads capture meaningful and prevalent attributes from the data. These components together enable CD to outperform existing methods in scenarios lacking class priors, as demonstrated by extensive experiments on benchmark datasets including CIFAR-10, CIFAR-100, STL-10, and ImageNet-10. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
@@ 1 Introduction
Recent contrastive learning methods [3, 4, 5] have incorporated class priors to enhance performance. By using class counts as prior knowledge during network construction, these approaches achieve state-of-the-art results in unsupervised clustering tasks, indicating that class information can lead to more discriminative feature representations. However, this reliance on class priors limits their applicability in real-world scenarios where class information is uncertain.
To address this limitation, we propose Contrastive Disentangling (CD), a novel framework that refines feature extraction and disentanglement in unsupervised learning without relying on class priors. CD integrates multi-level contrastive learning with a disentanglement loss to ensure feature diversity and independence, making it applicable to scenarios lacking class information. Our approach avoids complex additional modules such as pseudo-labeling, focusing instead on the effectiveness of the core framework.
@@
As shown in Figure 1, CD utilizes two levels of contrastive learning: instance-level and feature-level. Instance-level contrastive learning distinguishes image instances in the feature space, while feature-level contrastive learning ensures that distinct prediction heads capture diverse and complementary feature attributes. The feature-level module employs the LIME (Local Interpretable Model-Agnostic Explanations) technique [6] to visualize feature predictions, offering an intuitive interpretation of the model’s behavior. In summary, the main contributions of this paper are:
@@
    Feature Prediction Heads: We introduce feature prediction heads as a novel component in the feature learning process. Through the use of a normalized entropy loss, these heads effectively capture common features prevalent in the dataset, enhancing the representational capacity and generalization of the learned features.
@@
    Feature-Level Contrastive Learning: We propose a feature-level contrastive learning strategy to encourage different feature prediction heads to focus on distinct and complementary features, promoting effective feature disentanglement.
@@
    Simple and Effective Framework: We present a straightforward and effective framework that improves fine-grained feature learning in unsupervised tasks without relying on class priors. Extensive experiments on benchmark datasets demonstrate the superiority of CD, particularly in NMI scores. Ablation studies confirm the effectiveness of each component, providing new insights into unsupervised representation learning.
@@ 2.1 Unsupervised Representation Learning
Unsupervised representation learning seeks to extract meaningful features from unlabeled data. This area is dominated by two main approaches: generative models and contrastive learning. Generative models [7, 8, 9] aim to generate samples that resemble the true data distribution by modeling the underlying latent space. Despite their success in data generation, these models often struggle to learn invariant feature representations—features that remain consistent under various input transformations such as rotation, translation, and scaling [10].
To overcome these limitations, instance-level contrastive learning methods [2, 11, 12, 13] have been developed. These approaches enhance the model’s ability to extract invariant features by maximizing the similarity between different augmented views of the same instance while minimizing the similarity between different instances. As a result, contrastive learning has proven to be highly effective across various downstream tasks, offering robust performance in unsupervised learning.
@@ 2.2 Utilization of Class Priors
Several recent approaches [4, 5, 14] have incorporated class prior information to improve model performance by constraining the final outputs based on known class counts. These methods often achieve state-of-the-art results in unsupervised clustering tasks by guiding the model to learn more discriminative feature representations. Additionally, pseudo-labeling techniques are sometimes employed to iteratively refine cluster assignments, further boosting performance.
However, these methods are limited by their reliance on class priors, making them unsuitable for real-world tasks where class labels are often unknown or ambiguous. This reliance highlights the need for methods that can operate effectively without class information.
@@ 2.3 Multi-level Contrastive Learning
Multi-level contrastive learning has emerged as a promising direction within contrastive learning research. The introduction of cluster-level contrastive learning by Li et al. [4] for clustering tasks has opened new avenues for exploring more structured representations. However, methods that partition data based on predefined categories may miss the fine-grained features inherent in more complex datasets, limiting their ability to capture nuanced representations.
Building on this insight, we propose the Contrastive Disentangling (CD) framework, which leverages feature space representations as inputs to multiple feature prediction heads. By combining feature-level and instance-level contrastive learning, CD encourages a decoupling of distinct feature heads, enabling the model to learn more diverse and fine-grained feature representations. This approach advances unsupervised representation learning without relying on class priors, offering a more flexible and robust solution for a variety of tasks.
@@ 3.1 Proposed Contrastive Disentangling Framework
The proposed Contrastive Disentangling (CD) framework leverages a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
@@
    Feature Predictor: To capture diverse and disentangled feature attributes, the latent representations are further processed by a feature predictor $h(\cdot)$. This module, structurally similar to the instance projector but with a sigmoid activation in the final layer, outputs predictions $\mathbf{y}^{(1)}=h(\mathbf{z}^{(1)})$ and $\mathbf{y}^{(2)}=h(\mathbf{z}^{(2)})$. Each $\mathbf{y}^{(v)}$ is a matrix of size $N\times K$, where $K$ denotes the number of feature heads. Each feature head is designed to capture a distinct attribute of the data.
@@
    Loss Functions: The total loss function of CD integrates three components: the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, the feature-level contrastive loss $\mathcal{L}_{\text{feat}}$, and the normalized entropy loss $\mathcal{L}_{\text{entropy}}$. These losses work synergistically to promote feature diversity, encourage independence among the learned features, and prevent feature collapse.
@@ 3.2 Instance-level Contrastive Learning
Instance-level contrastive learning aims to learn discriminative features by bringing representations of different augmented views of the same instance closer while pushing apart representations of different instances. This dual objective enhances the model’s ability to distinguish between distinct instances and recognize invariant features across different augmentations.
@@
for $i,j\in\{1,2,\dots,2N\}$. The purpose of using cosine similarity is to measure how closely related two feature vectors are, with a higher value indicating greater similarity. Building on this similarity measure, we define the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, which aims to maximize the similarity between positive pairs (i.e., the other augmented view of the same instance) and minimize the similarity between negative pairs (i.e., views of different instances). For each sample $i$, the instance-level contrastive loss is formulated as:
@@
where $\text{pos}(i)$ denotes the index of the positive sample corresponding to $i$, $\tau_{\text{inst}}$ is a temperature parameter controlling the concentration level of the distribution, and $\mathbbm{1}_{[k\neq i]}$ is an indicator function that equals 1 when $k\neq i$ and 0 otherwise. This formulation follows the NT-Xent loss used in previous work [2, 17, 18, 19]. The total instance-level contrastive loss is computed by averaging over all $2N$ samples:
@@ 3.3 Disentanglement Loss
To ensure that the learned feature representations are both diverse and independent, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes diversity by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the captured features are meaningful and prevalent across the dataset.
@@ 3.3.1 Feature-Level Contrastive Loss
The feature predictor $h(\cdot)$ produces predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads. To apply feature-level contrastive learning, we concatenate the transposed outputs of the two views:
@@
The feature-level contrastive loss encourages each feature head to capture distinct attributes by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads. For feature head $i$, the loss is defined as:
@@ 3.3.3 Combined Disentanglement Loss
where $\alpha$ is a hyperparameter that controls the balance between encouraging diversity and ensuring meaningful feature representations. By jointly optimizing these two objectives, the feature heads are encouraged to capture both diverse and informative patterns, leading to higher-quality feature disentanglement.
@@ 3.4 Total Loss
By optimizing both the instance and feature-level objectives, our framework achieves a robust balance between learning discriminative instance-level representations and promoting diverse, meaningful feature representations without relying on class priors. Algorithm 1 summarizes the proposed method.
@@ 4.1 Datasets
We conducted experiments on four widely-used datasets: CIFAR-10, CIFAR-100 [20], STL-10 [21], and ImageNet-10 [22]. These datasets serve as benchmarks to rigorously evaluate the effectiveness of our proposed Contrastive Disentangling (CD) framework. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes.
@@ 4.4 Experimental Results
Dataset Metrics K-means AE DeCNN DAE VAE DCGAN CC* CD-128 CD-256
@@
[…]
Final output. Table 3 compares CD to models that use class number information. CD demonstrated strong performance across all metrics, particularly on more complex datasets like STL-10 and ImageNet-10. While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is expected as CD does not use class number information. Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data.
@@
Dataset Metrics JULE* DEC* DAC* DCCM* PICA* CC* CD-128 CD-256
[…]
@@ 4.5.3 Feature Prediction Head
We evaluated the effect of removing the feature prediction head, which reduces the model to an instance-level contrastive learning framework similar to SimCLR. As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation, highlighting the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results).
In comparison to SimCLR under identical training settings, our model demonstrates significant improvements: on the STL-10 dataset, NMI increases by 13.4%, ARI by 23.1%, and ACC by 10.9%; on ImageNet-10, NMI improves by 26.8%, ARI by 110.2%, and ACC by 53.1%. These results present that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, leading to improved clustering and representation quality in the final output.

### v3

@@ Abstract
Recent advances in unsupervised representation learning often rely on knowing the number of classes to improve feature extraction and clustering. However, this assumption raises an important question: is the number of classes always necessary, and do class labels fully capture the fine-grained features within the data? In this paper, we propose Contrastive Disentangling (CD), a framework designed to learn representations without relying on class priors. CD leverages a multi-level contrastive learning strategy, integrating instance-level and feature-level contrastive losses with a normalized entropy loss to capture semantically rich and fine-grained representations. Specifically, (1) the instance-level contrastive loss separates feature representations across samples; (2) the feature-level contrastive loss promotes independence among feature heads; and (3) the normalized entropy loss ensures feature diversity and prevents feature collapse. Extensive experiments on CIFAR-10, CIFAR-100, STL-10, and ImageNet-10 demonstrate that CD outperforms existing methods in scenarios where class information is unavailable or ambiguous. The code is available at https://github.com/Hoper-J/Contrastive-Disentangling.
@@ 1 Introduction
Although recent contrastive learning methods [3, 4, 5] have leveraged class priors to improve performance, their reliance on predefined class counts limits their applicability in real-world scenarios where class information may be uncertain or unavailable. This dependence not only restricts the adaptability of these models but also hampers their ability to generalize to diverse tasks that lack explicit class labels.
To address this limitation, we propose Contrastive Disentangling (CD), a novel unsupervised learning framework that improves feature extraction and disentanglement without requiring class priors.
@@
Our approach stems from a detailed observation of the model’s outputs. As illustrated in Figure 1, the rows of the instance projector correspond to the latent representations of the samples, while the columns of the feature predictor represent the predictions of fine-grained features across the entire batch. Traditional contrastive learning methods focus predominantly on instance-level feature representations, yet we raise a new question: can the focus be shifted to the features themselves? Motivated by this insight, we extend contrastive learning to the feature level, treating each feature prediction head as an independent “feature instance” to better capture the underlying attributes of the data. This conceptual shift forms the basis of our CD framework, which integrates multi-level contrastive learning with disentanglement loss to ensure feature diversity and independence, while maintaining a simple and efficient architecture. In summary, the main contributions of this paper are:
@@
    We introduce feature prediction heads and a novel feature-level contrastive approach, enabling the model to capture finer-grained features within a simple and effective framework, providing a fresh and generalizable perspective on contrastive learning.
@@
    We incorporate a normalized entropy loss to complement feature-level contrastive learning, ensuring that the model captures prevalent and diverse features across the dataset.
@@
    To the best of our knowledge, this is the first contrastive learning method that achieves semantically rich final outputs without relying on class priors. By defining the model architecture independently of class information, our approach allows researchers to intuitively identify images with similar fine-grained features by examining the final model outputs, which exhibit the same level of semantic richness as the feature extraction layer. Moreover, the model offers researchers the flexibility to adjust the number of feature heads, enabling control over the granularity of the features learned.
Our method demonstrates superior feature extraction capabilities compared to existing unsupervised frameworks and remains highly competitive against methods that rely on class information to constrain model outputs. Additionally, under identical training settings, the final outputs of CD exhibit richer semantics than SimCLR. This improvement is particularly evident in clustering tasks. For example, in the STL-10 dataset, CD achieves a 13.4% increase in NMI, a 23.1% increase in ARI, and a 10.9% improvement in ACC. On the ImageNet-10 dataset, CD shows even greater enhancements, with a 26.8% increase in NMI, a 110.2% improvement in ARI, and a 53.1% increase in ACC.
@@ 2.1 Unsupervised Representation Learning
Unsupervised representation learning focuses on extracting meaningful features from unlabeled data, with two primary approaches: generative models and contrastive learning. Generative models [6, 7, 8] aim to generate samples that resemble the true data distribution by modeling the underlying latent space. Despite their success in data generation, these models often struggle to learn invariant feature representations—features that remain consistent under various input transformations such as rotation, translation, and scaling [9].
In contrast, instance-level contrastive learning methods [2, 10, 11, 12] improve the model’s ability to extract invariant features by maximizing the similarity between augmented views of the same instance while minimizing the similarity between different instances. This strategy has consistently demonstrated strong performance across various downstream tasks, positioning contrastive learning as a highly effective method in unsupervised learning.
@@ 2.2 Utilization of Class Priors
Recent methods [4, 5, 13] incorporate class priors to improve performance by constraining the learning process with known class counts and refining cluster assignments through pseudo-labeling. These approaches have yielded state-of-the-art results in unsupervised clustering, as class priors guide the model towards more discriminative representations.
However, such reliance on class priors significantly limits the adaptability of these models, especially in real-world scenarios where class information is incomplete, ambiguous, or unavailable. This motivates the need for approaches that can effectively capture meaningful and discriminative representations without relying on predefined class information.
@@ 2.3 Multi-level Contrastive Learning
Multi-level contrastive learning has recently emerged as a novel direction within contrastive learning research. For example, cluster-level contrastive learning by Li et al. [4] was introduced for clustering tasks, opening new avenues for structured representation learning. However, methods that partition data based on predefined categories may overlook fine-grained features in more complex datasets, limiting their ability to capture subtle and nuanced patterns.
Building on these observations, our framework extends contrastive learning to both the instance and feature levels. By treating each feature prediction head as an independent “feature instance,” our approach encourages each head to capture distinct, complementary attributes, which is a significant departure from traditional instance-level contrastive learning. This decoupling of feature heads allows CD to learn diverse and fine-grained representations, making it a flexible and powerful method for unsupervised representation learning without relying on class priors.
@@ 3.1 Proposed Contrastive Disentangling Framework
The proposed framework employs a multi-level contrastive learning strategy to achieve fine-grained representation learning without the need for class priors, as illustrated in Figure 2.
@@
    Feature Predictor: The latent representations are processed by the feature predictor $h(\cdot)$ to capture diverse and disentangled attributes. This module, structurally similar to the instance projector but with a sigmoid activation in the final layer, outputs predictions $\mathbf{y}^{(1)}=h(\mathbf{z}^{(1)})$ and $\mathbf{y}^{(2)}=h(\mathbf{z}^{(2)})$. Each $\mathbf{y}^{(v)}$ is a matrix of size $N\times K$, where $K$ denotes the number of feature heads. Each feature head is designed to capture a distinct attribute of the data.
@@
    Loss Functions: The total loss function of CD integrates three components: the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, the feature-level contrastive loss $\mathcal{L}_{\text{feat}}$, and the normalized entropy loss $\mathcal{L}_{\text{entropy}}$. These losses work together to ensure that the learned features are diverse, independent, and well-distributed, while avoiding feature collapse.
@@ 3.2 Instance-level Contrastive Learning
Instance-level contrastive learning seeks to learn discriminative representations by bringing augmented views of the same instance closer while pushing apart representations of different instances. This dual objective improves the model’s ability to differentiate between instances and capture invariant features across augmentations.
@@
for $i,j\in\{1,2,\dots,2N\}$. Cosine similarity measures the relationship between latent representations, with higher values indicating greater similarity. Building on this similarity measure, we define the instance-level contrastive loss $\mathcal{L}_{\text{inst}}$, which aims to maximize the similarity between positive pairs (i.e., the other augmented view of the same instance) and minimize the similarity between negative pairs (i.e., views of different instances). For each sample $i$, the instance-level contrastive loss is formulated as:
@@
where $\text{pos}(i)$ denotes the index of the positive sample corresponding to $i$, $\tau_{\text{inst}}$ is a temperature parameter controlling the concentration level of the distribution, and $\mathbbm{1}_{[k\neq i]}$ is an indicator function that equals 1 when $k\neq i$ and 0 otherwise. The total instance-level contrastive loss is computed by averaging over all $2N$ samples:
@@ 3.3 Disentanglement Loss
To ensure that the predicted feature attributes capture meaningful variation across the dataset, we introduce two complementary loss functions: the feature-level contrastive loss, which promotes independence by encouraging each feature head to focus on distinct attributes, and the normalized entropy loss, which ensures that the learned features are well-distributed and representative of the overall data distribution.
@@ 3.3.1 Feature-Level Contrastive Loss
The feature predictor $h(\cdot)$ generates predictions $\mathbf{y}^{(1)}$ and $\mathbf{y}^{(2)}$, each of size $N\times K$, where $K$ is the number of feature heads. In our framework, each feature head is treated as a distinct “feature instance” that captures different characteristics of the dataset. By introducing feature-level contrastive learning, our method uniquely enables each head to capture complementary and independent attributes. To maintain consistency in the formulation, we transpose the output matrix of the feature heads and concatenate the outputs from both views:
@@
The feature-level contrastive loss promotes the independence of feature heads by maximizing the similarity between corresponding heads across different views while minimizing the similarity between different heads. For feature head $i$, the loss is defined as:
@@ 3.3.3 Combined Disentanglement Loss
where $\alpha$ is a hyperparameter that controls the balance between encouraging diversity and ensuring meaningful feature representations. By simultaneously optimizing these two objectives, the feature heads are encouraged to capture both diverse patterns and informative representations, leading to better feature disentanglement.
@@ 3.4 Total Loss
By jointly optimizing instance and feature-level objectives, our framework achieves a unique balance between learning discriminative representations and ensuring feature diversity, all without relying on class priors. This makes our approach especially suited for complex unsupervised learning tasks where class information is scarce or completely unavailable, providing a robust solution for learning fine-grained and diverse representations. Algorithm 1 summarizes the proposed method.
@@ 4.1 Datasets
We conducted experiments on four widely-used datasets: CIFAR-10, CIFAR-100 [16], STL-10 [17], and ImageNet-10 [18]. These datasets serve as benchmarks to rigorously evaluate the effectiveness of our approach. Table 1 provides an overview of the dataset splits, sample sizes, and the number of classes.
@@ 4.4 Experimental Results
Dataset CIFAR-10 CIFAR-100-20 STL-10 ImageNet-10
Metrics NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC
K-means 0.087 0.049 0.229 0.084 0.028 0.130 0.125 0.061 0.192 0.119 0.057 0.241
AE 0.239 0.169 0.314 0.100 0.048 0.165 0.250 0.161 0.303 0.210 0.152 0.317
DeCNN 0.240 0.174 0.282 0.092 0.038 0.133 0.227 0.162 0.299 0.186 0.142 0.313
DAE 0.251 0.163 0.297 0.111 0.046 0.151 0.224 0.152 0.302 0.206 0.138 0.304
VAE 0.245 0.167 0.291 0.108 0.040 0.152 0.200 0.146 0.282 0.193 0.168 0.334
GAN 0.265 0.176 0.315 0.120 0.045 0.151 0.210 0.139 0.298 0.225 0.157 0.346
DCGAN 0.265 0.176 0.315 0.120 0.045 0.151 0.210 0.139 0.298 0.225 0.157 0.346
CC* 0.670 0.523 0.684 0.433 0.244 0.431 0.608 0.443 0.594 0.800 0.695 0.846
@@
Final output. Table 3 compares CD to models that use class number information. CD demonstrated strong performance across all metrics, particularly on more challenging datasets like STL-10 and ImageNet-10. While ARI and ACC were slightly lower on CIFAR-10 and CIFAR-100 compared to CC, this is primarily due to the lower complexity of these smaller datasets, which limits the advantage of CD’s ability to disentangle fine-grained features. In datasets with more intricate patterns and greater feature diversity, CD’s strengths are more effectively utilized. Importantly, CD achieved an ACC of 0.734 with a batch size of 256, surpassing PICA (0.713) despite not using additional unlabeled data, which consists of 100,000 samples.
@@
Dataset CIFAR-10 CIFAR-100-20 STL-10 ImageNet-10
Metrics NMI ARI ACC NMI ARI ACC NMI ARI ACC NMI ARI ACC
JULE* 0.192 0.138 0.272 0.103 0.033 0.137 0.182 0.164 0.277 0.175 0.138 0.300
DEC* 0.257 0.161 0.301 0.136 0.050 0.185 0.276 0.186 0.359 0.282 0.203 0.381
DAC* 0.396 0.306 0.522 0.185 0.088 0.238 0.366 0.257 0.470 0.394 0.302 0.527
DCCM* 0.496 0.408 0.623 0.285 0.173 0.327 0.376 0.262 0.482 0.608 0.555 0.710
PICA* 0.591 0.512 0.696 0.310 0.171 0.337 0.611 0.531 0.713 0.802 0.761 0.870
CC* 0.705 0.637 0.790 0.417 0.221 0.421 0.622 0.539 0.670 0.859 0.822 0.893
@@ 4.5.3 Feature Prediction Head
We evaluated the effect of removing the feature prediction head, which reduces the model to an instance-level contrastive learning framework similar to SimCLR. As shown in Table 6, the feature prediction head significantly boosts performance, and its removal results in substantial degradation. This highlights the critical role of feature-level contrastive learning in effective disentanglement (see Table 9 for backbone results). These findings demonstrate that our feature-level contrastive learning approach enables the model to capture more distinct and diverse semantic features, resulting in improved representation quality in the final output.